What is the 5th root of a^5b^10?
Found 2 solutions by richwmiller, dabanfield:
Answer by richwmiller(17219)

fifth root is written ^(1/5)
(a^5b^10)^(1/5)
(a^5)^(1/5)*(b^10)^(1/5)
(a^(5/5)*(b^(10/5)
a*b^2
because 5/5=1 and 10/5=2

Answer by dabanfield(803)

What is the 5th root of a^5b^10?
The fifth root of a^5*b^10 = [fifth root of a^5 * fifth root of b^10]
Fifth root of a^5 = a^(1/5*5) = a^1 = a
Fifth root of b^10 = b^(1/5*10) = b^(10/5) = b^2
So we have a*b^2 for our answer.